A boat is moving due east in a region where the earth’s magnetic field is 5 11 5.0 10 NA m - -- × due north and
horizontal. The boat carries a vertical aerial 2m long. If the speed of the boat is 1.50 ms-1
, the
magnitude of the induced emf in the wire of aerial is :
?

Grade:11

1 Answers

Kartikeya
6 years ago
Hey mates,We know emf induced= b×v×l = 5×10^-5 × 1.50×2 ✓✓✓✓=10×10^-5×1.50=15×10^-5I hope you like this answer, if you did, then stay Sexy babe#PeaceOut